1. A sample of fresh banana is dried to remove all water. The dried banana weighs 3.0g. This 3.0g sample is burned and the final ash weight is 1.0g. If 83% of fresh banana tissue is water:

a. What was teh original wet weight of the 3.0 gram sample of dried banana? Show your work.

b. When the dried sample of banana was burned, what became of the organic molecules that made up the banana tissue?

c. When the dried sample of banana was burned, what became of the inorganic compounds in the banana tissue?

d. What is the organic content of banana tissue expressed as a percentage of fresh weight? Show your work.

e. What is the inorganic content expressed as a precentage of fresh weight? Show your work.

f. In lab, burning 3.0g of dry tissue will not be practically possible because 3.0g is too large a mass to fit safely inside the crucible. So you wisely decide to burn only 1.0g of the 3.0g total dried banana. Based on the known composition of banana tissue, how many grams of ash do you expect to find in the crucible? Show your work.
Answer #1

a. Given that, 83% of fresh banana tissue is water (i.e. 17% of fresh banana tissue is dry matter) and the dried banana weighs 3.0g.

17 g of dry banana is obtained from 100 g fresh banana

3 g of dry banana will be obtained from (100/17)×3 = 17.65 g of fresh banana sample

b. The organic molecules underwent oxidation to form gases and were lost to the environment.

c. The inorganic molecules were recovered as ash.

d. Organic content of dry banana = (3 - 1) = 2 g

Fresh weight = 17.65 g

% organic content = 2/17.65 × 100% = 11.33%

e. Similarly, inorganic content = 1/17.65 × 100% = 5.67%

f. 3 g dried banana gave 1 g ash

1 g dried banana will give (1/3) = 0.33 g ash